Pol He3 Target Meeting June 1, 2020

From HallCWiki
Jump to: navigation, search

Members present: Jian-ping, Brad, Junhao, Mingyu, Melanie, Murchhana, Arun, Todd, and Christopher

Begin time: 1:30 p.m. End time: 2:54 p.m.

Junhao

  • Laser control software was upgraded; need to redo EPICS integration and test it with the new lasers once the lab reopens
  • Ordered the spectrometer for the transverse laser; two weeks lead time
  • Ordered RF current monitor (4-5 weeks lead time)
  • Need to ship back a malfunctioning laser to Raytum once JLab reopens
    • ask about a return authorization number
  • Dutch AFP loss: most testing was done during commissioning
    • There are only two data files for the standard correction field settings
    • The AFP loss gets larger for longer wait times (DS up and down), perhaps due to the difference in polarization loss and diffusion between the different chambers
    • It's suspicious that the target chambers (DS up and down, UP up and down) have higher loss than the pumping chamber (PC up and down)
    • Check the 9.20% “DS down” value!
    • It's important to make sure the convection is stable when doing AFP loss studies (must know the conditions before and after measurements)
    • Need to combine this info with the temperature tests

Mingyu

  • Worked on obtaining average beam current for target depolarization studies (want to ultimately build a model to more accurately know the target polarization during a certain run)
    • Used "ibcm1" from myStats Mean value for archived ibcm1 variables. This includes any beam on and beam off time.
    • Suggestion by Brad: Stripchart the ibcm1 value, look for where the current is stable and where it overlaps with two polarimetry measurements
  • Remaining work:
    • Missing info for Dutch: target polarization test just before it was installed (2/10)
    • With Junhao’s polarimetry analysis, correct the NMR calibration constants and put corrected target polarimetry results into the database

Murchhana

  • Started a document for the field direction measurements with updated errors

Melanie

  • Revised error propagation in EPR analyses (fixed bug in script and applied it to all calibrations)
  • Will add Junhao's updated AFP loss values

Chris

  • Finished testing Butterball and Tommy, need to pack them up and shift to JLab
  • Nothing going on with making a new cell until the pandemic shutdown is lifted

JP

  • MEDCON5 is expected to be lifted sometime next week. Then experiment running will begin 4-6 weeks later
  • Sent an email to Steve Wood to talk about the target tasks and who will be onsite once the lab reopens. A copy of its contents is as follows:
    • Task List for re-starting the polarized He3 target operation in Hall C (assuming we already have documents re-approved before starting up):
      • First week: Start-up (Junhao, Mingyu, Bill, Brad, an Arun)
        • Check and restart the control system/instrumentation in the counting house and in the hall (including power supplies, motion system, laser in the laser room only (laser in the hall needs to be after optics clean-up) (Junhao/Mingyu/Bill/Arun/JP)
        • Open up the enclosure *two panels) (JP/Arun)
        • Clean up the optics (Junhao/Mingyu)
        • Put Hall in laser controlled access (JP/Arun)
        • Check laser alignment, laser polarization (JP/Junhao/Mingyu)
        • Check EPR setup inside the enclosure and laser spectro-analyzer, laser wave-plates remote control (Junhao/Mingyu)
        • Close up the enclosure and put Hall out of laser controlled access (JP/Arun)
        • Start up optical pumping, testing NMR/EPR/pNMR systems (Junhao/Mingyu/JP)
        • Test new Kepco power supply (Bill/Junhao)
      • Second week: (same people plus Todd, if feasible, plus the two students, Melanie and Murchhana)
        • Spin-up and making sure the target works in all 4 polarization directions
        • EPR calibrations, AFP loss study, temperature study for all configurations
        • Starting optimization for all 4 configurations
        • Checkout reference cell system and cooling jets
      • Third week: (Should have Melanie and Murchhana on board to perform polarimetry study)
        • Continue optimization of target performance in all 4 configurations
        • Prepare one new cell